40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 29 14:22:03 2011, Zach, Update, RF System, EOM temp stabilization performance Screen_shot_2011-11-29_at_1.33.50_PM.pngScreen_shot_2011-11-29_at_1.34.06_PM.pngScreen_shot_2011-11-29_at_1.41.03_PM.png
    Reply  Tue Nov 29 15:25:29 2011, steve, Update, RF System, EOM temp stabilization performance 2days.png
       Reply  Tue Nov 29 18:17:27 2011, Zach, Update, RF System, EOM temp stabilization performance 
    Reply  Tue Nov 29 22:10:18 2011, kiwamu, Update, RF System, RFAM fluctuation reduced RFAM_withEOMheater_edit.png
       Reply  Tue Nov 29 23:03:34 2011, Zach, Update, RF System, RFAM fluctuation reduced 
          Reply  Wed Nov 30 03:01:55 2011, kiwamu, Update, RF System, RFAM fluctuation reduced 
Message ID: 6035     Entry time: Tue Nov 29 14:22:03 2011     Reply to this: 6036   6044
Author: Zach 
Type: Update 
Category: RF System 
Subject: EOM temp stabilization performance 

I left the EOM stabilization running overnight, so we can finally see how the EOM temperature stabilization does over long periods of time.

Here are both long-term (~13-hr) and short-term (1-hr) trends of the EOM temperature and the heater drive level. From the long trend you can see that the heater departed the steady value of ~7.8V that I observed last night to accommodate the diurnal heating of the lab in the morning---the temperature was held near zero offset.

From the short-term trend, there are 2 things to notice:

  1. We are still very close to the digitization noise level for both signals. This is bad, because we want to look at the residual noise level, etc.
  2. There appears to be some strange sort of disturbance of f~0.01-0.02 Hz. I'm not sure what causes the strange shape

Screen_shot_2011-11-29_at_1.33.50_PM.png Screen_shot_2011-11-29_at_1.34.06_PM.png

 

Finally, here is a trend over the last ~24 hours of the EOM temperature, heater drive level, and the 11- and 55-MHz Stochmon signals. I believe that the abrupt shelves noticeable on the Stochmon trends are when c1sus was turned on and shut down, respectively (I'm not sure why that causes the signals to die, but the times seem right, and nothing obvious happens to the EOM temp stabilization signals at either time). The controller was turned on at ~8:40 UTC, and you can see that the Stochmon signals quiet down a lot right at that time. There is some residual drift (common-mode to both RF frequencies), which is most likely caused by a drift in some other parameter (e.g. laser frequency or power).

Screen_shot_2011-11-29_at_1.41.03_PM.png

I took some relatively inconclusive power spectra and coherence measurements, but I'd rather wait until we have an uncontrolled data stretch with which to compare. I think what we should do now is disconnect the controller and then let it sit for a while.

 

ELOG V3.1.3-